carmelo is 12 yrs older than Arwin.One-third of Carmelo's age is equal to one-half of Arwin's age. how old is each now?
Answer by stanbon(75887)

carmelo is 12 yrs older than Arwin.One-third of Carmelo's age is equal to one-half of Arwin's age. how old is each now?
--------
Equations:
c = a + 12
(1/3)c = (1/2)a
----
Modify:
c = a + 12
2c = 3a
----
Substitute for "c" and solve for "a":
2(a+12) = 3a
2a + 24 = 3a
a = 24 (Arwin's age)
Solve for "c":
c = a + 12
c = 24+12 = 36 (Carmelo's age)
=================
Cheers,
Stan H.
=================